OpenStudy (anonymous):

(4c^-3d)0/25(d+3)0 answer with a positive exponent step by step please

jimthompson5910 (jim_thompson5910):

Those 0's are exponents right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Anything with an exponent of 0 is 1 So the numerator becomes 1 and the expression (d+3)^0 also turns into 1

OpenStudy (anonymous):

1/25?

OpenStudy (anonymous):

how did you get 1/25? :/?

jimthompson5910 (jim_thompson5910):

everything else basically turns to 1 which cancel out

OpenStudy (anonymous):

If the 25 isn't to the power of 0 that is.

OpenStudy (anonymous):

so sense the 0 power is there it makes everything cancel out?

OpenStudy (anonymous):

(4c^-3d)^0 = 1 (d+3)^0 = 1 (4c^-3d)^0/25(d+3)^0 = 1/25*1 = 1/25

OpenStudy (anonymous):

ahh ok gotcha thanks